OpenStudy (anonymous):

Simplify square root of 13 over 11 .

OpenStudy (anonymous):

I take that to mean \[\sqrt{13/11} = \sqrt{13}/\sqrt{11}\] Again we rationalize by multiplying sqrt11/sqrt11 \[(\sqrt{13}/\sqrt{11})*(\sqrt{11}/\sqrt{11})\] and we get \[\sqrt{143}/11\]

OpenStudy (anonymous):

That was exactly it , thank you so much .